OpenStudy (yadu123):

Aminocyclohexane has a percent composition of 64% carbon, 11.5% hydrogen and 24.5% nitrogen. 8 moles of it weighs 900 grams. What is the compounds empirical formula? What is this compounds molecular formula? How many hydrogen atoms are in a 3 mole sample?

OpenStudy (yadu123):

@Vocaloid

OpenStudy (yadu123):

@paki

OpenStudy (anonymous):

If 8 mol aminocyclohexane is 900 g minocyclohexane , then the molar mass is 900 g/8 mol = 112.5 g/mol. Assume there's 100 g of aminocyclohexane, then you can say in that sample there's 64 g C, 11.5 g H, and 24.5 N. Use the molar mass to convert these masses to moles. (64 g C)(1 mol C/12 g C) = 5.3 mol C (11.5 g H)(1 mol H/1 mol H) = 11.5 mol H (24.5 mol N)(1 mol N/28 g N) = 0.875 mol N Nitrogen has the smallest amount so divide all the amounts by 0.875 to get the subscripts C: 5.3/0.875 = 6 H: 11.5/0.875 = 13 N: 0.875/0.875 = 1 So the empirical formula is \(C_6H_{13}N\).

OpenStudy (yadu123):

Thx

OpenStudy (yadu123):

Thx